A crystalline salt Na2SO4.xH2O on heating losses 55.9 per cent of its mass. The formula of crystalline salt is---

Grade:12

1 Answers

Manas Satish Bedmutha
9 years ago
Mol. wt. of hydrated salt = 46 + 96 + 18x = 142 +18x
Wt. of water lost = 18x
So, by given, 18x / (142+18x) = 55.9
On sloving, we get, x=10
So Formula = Na2SO4. 10 H2O